the length of the 2 shorter sides of a right angleare 20 and 15.another right angle of equal area has its shortside 10cm long.what is the length of the side perpendiculars to this?

Answer by stanbon(75887)

The length of the 2 shorter sides of a right angle are 20 and 15.another right angle of equal area has its shortside 10cm long.what is the length of the side perpendiculars to this?
---------------------
The triangles are similar.
Area of the 1st triangle = (1/2)(20)(15) = 150 sq units
---------
Let the length of the unknown side of the 2nd right triangle be "x".
Area = (1/2)10x
150 = 5x
x = 30 units (lenth of the unknown side)
=============================================
Cheers,
Stan H.